OpenStudy (anonymous):

3y^3-6y^2-3y+6 factor completely. pleaseeee

jhonyy9 (jhonyy9):

\[3y(y ^{2}-1)-6(y ^{2}-1)=(3y-6)(y ^{2}-1)=3(y-2)(y ^{2}-1)\]